# SeNeDiF-OOD: Semantic Nested Dichotomy Fusion for Out-of-Distribution Detection Methodology in Open-World Classification. A Case Study on Monument Style Classification

### v2

@@ 1 Introduction
Drawing a parallel with multi-class classification, where each class would represent either an in-distribution (ID) input or a type of OOD input, we can consider nested dichotomies techniques [15] to address this problem. Nested dichotomies break down classification tasks into multiple binary subproblems, each constrasting one category against the rest. During testing, each step determines whether the sample belongs to the split class. If it does not, the process continues to the next dichotomy. This sequence repeats until the dichotomy that assigns a class to the input is reached. Generally, nested dichotomies are less favored over other decomposition techniques like one-vs-one or one-vs-all [28] because, despite being more efficient, their performance depends significantly on the hierarchy of divisions within the dichotomies. However, when categories are already hierarchically structured, this limitation turns into an advantage.
@@
Many computer vision deployments still operationalize OOD recognition as a single scalar score that separates ID from OOD under an essentially unstructured notion of novelty. This abstraction is misaligned with open-world, user-facing pipelines, where OOD inputs arise from heterogeneous mechanisms and demand different responses, ranging from unrelated content to semantically adjacent but unsupported classes and acquisition degradations. Consequently, a monolithic score offers limited interpretability and provides weak guidance for system-level handling. A missing ingredient is a principled methodology that decomposes OOD recognition into semantically meaningful decisions, enabling robust filtering while characterizing distinct OOD modes.
@@
A structured view of novelty can be obtained by organizing the input space into successive semantic checkpoints rather than treating it as a flat ID versus OOD partition. In such a design, each checkpoint captures a coarse-to-fine distinction that mirrors how a practitioner would rule out implausible content before committing to a specific interpretation. This hierarchical organization encourages the use of specialized criteria at each stage and constrains error propagation, since ambiguous samples are intercepted early by broad semantic tests instead of being forced into a final decision. Beyond improving reliability, this methodology naturally yields a diagnostic outcome that indicates where and why an input becomes incompatible with the intended operating domain, which is valuable for monitoring, curation, and data acquisition strategies.
@@
In this paper, we introduce SeNeDiF-OOD (Semantic Nested Dichotomies Fusion for OOD detection), a new methodology that casts OOD recognition as a structured fusion process over a semantic cascade of dichotomous decisions. In particular, SeNeDiF-OOD combines hierarchical decision-level fusion with semantic priors embedded in the nested dichotomies topology. These priors provide human-aligned guidance that facilitates OOD detection. At each level of the hierarchy, the method also integrates heterogeneous expert fusion, leveraging the representational capabilities of AI models to better separate the underlying dichotomies. Overall, this design enables a principled fusion of human knowledge and machine intelligence, which can be beneficial for facilitating OOD detection, a setting in which many computer vision models still struggle [61].
@@
Using a dataset composed of images uploaded by MonuMAI users, complemented with additional images obtained from public datasets covering different domains, we conducted an experiment to compare SeNeDiF-OOD methodology against the MonuMAI deep learning model itself. The results confirm that SeNeDiF-OOD significantly reduces false detections of monuments compared to the original model while preserving ID images. Additionally, we analyze the system layer by layer, evaluating the effectiveness of each layer individually and drawing conclusions about the types of OOD images filtered at each stage.
@@
The structure of this paper is as follows. Section 2 provides the necessary background on out-of-distribution OOD detection and nested dichotomies. Section 3 establishes the theoretical foundations of the proposed SeNeDiF-OOD methodology, formulating a generalizable methodology for hierarchical decision fusion. Section 4 details the practical implementation of the SeNeDiF-OOD methodology within the MonuMAI framework, outlining the established layers and the training procedures for each. Section 5 presents the experiments conducted on this dichotomy-based system, comparing its performance with the original MonuMAI version and analyzing the results both globally and at the layer level. Finally, Section 6 concludes with a summary of the work and key takeaways.
@@ 2.1 Out-of-distribution detection
Out-of-distribution detection has emerged as a fundamental component for ensuring the robustness, safety, and trustworthiness of modern machine learning (ML) systems [55, 18]. In real-world deployment, models frequently encounter inputs that deviate significantly from the statistical distribution of their training data. These inputs, referred to as OOD samples, can originate from different semantic categories, domains, acquisition conditions, or even adversarial manipulations. If not properly identified, OOD samples may be misclassified with high confidence, potentially leading to severe consequences in high-stakes domains such as autonomous driving, healthcare, or financial systems.
The importance of OOD detection lies in its role as a safeguard mechanism. For example, an autonomous driving system trained primarily on urban street scenes should raise an alert when faced with unexpected rural or adverse weather conditions, instead of producing overconfident predictions that may compromise safety [23, 39]. Similarly, in medical imaging, diagnostic systems must detect when a sample corresponds to a rare disease or a novel variant not represented in training data, in order to avoid erroneous diagnoses [27, 53, 5]. Likewise, in natural language processing, the rise of foundation models and their applicability as autonomous agents makes it necessary to further develop robust techniques for handling text inputs that fall outside the scope anticipated during system design [26, 56]. The primary goals of OOD detection are therefore to improve system reliability by preventing harmful misclassifications, to facilitate safer deployment of ML models, and to enable iterative model refinement by flagging unexpected cases for human inspection [4].
@@ 2.2 Nested dichotomies for multi-class problems
Nested dichotomies [15] constitute a structured decomposition framework for tackling multi-class classification problems by reducing them into a hierarchy of binary decisions. Unlike traditional strategies such as one-vs-one or one-vs-all, nested dichotomies organize class discrimination into a binary tree structure, where each internal node corresponds to a binay classifier that separates a subset of classes into two disjoint partitions. This recursive partitioning continues until each node corresponds to a single class.
@@
3 The SeNeDiF-OOD methodology for image clasification
@@
Section 3.1 introduces a taxonomy of OOD image types, illustrating the variety of OOD scenarios that can arise across different applications. Section 3.2 formalizes the SeNeDiF-OOD architecture, constructing a semantically grounded hierarchy that maps specific OOD categories to distinct decision fusion nodes. Lastly, Section 3.3 situates the proposed methodology within the broader context of responsible artificial intelligence (AI), emphasizing key considerations such as robustness, explainability, and human alignment.
@@ 3.1 Open-world image classification: identifying OOD categories
  • Malicious OOD inputs. They are often stemming from adversarial attacks or other exploits that leverage the model weaknesses. These images are not only out-of-distribution but intentionally crafted to deceive the model, exploiting weaknesses in the decision boundaries of the models. Unlike accidental OOD inputs, these are targeted and present an elevated risk in high-stakes applications.
@@
  • Degraded images. These include inputs with heavy noise, extreme lighting conditions, motion blur or compression artifacts. While these images may still be in scope in terms of subject matter, the training distribution of the models is not likely covering these distortions comprehensively.
@@ 3.2 SeNeDiF-OOD world hierarchy and layers
  • The known-world: the exact training distribution. Finally, the known-world represents the precise training distribution on which the model was trained on. The last layer has the goal of filtering out the novel classes that are not represented in the training set, but are considered to belong to the full-problem world. For the last layer, it will be usually necessary to resort to a specialized OOD detector, since it may not be feasible to have a full distribution of the problem unknown world. Detectors with outlier exposure [22] may help to define a clearer border between known and unknown worlds and facilitate novelty detection. This world represents the comfort zone of the model, where it can make confident predictions based on its learned parameters. This world is the ID world: no OOD inputs are expected in this world.
This hierarchical structure of OOD categories can be naturally addressed using nested dichotomies, by directly extending the approach typically employed for multi-class problems. In this context, each OOD category can be treated as a distinct class, after excluding its internal categories to prevent overlap. Furthermore, the hierarchical organization of the categories enables the construction of a canonical and semantically meaningful dichotomy tree. At each node, the objective is to distinguish the outermost category from the remaining ones, progressively refining the classification until reaching the ID category in the final split. This design avoids reliance on a randomly generated dichotomy tree, providing a more structured and interpretable decision process.
@@
Figure 3: Schematic of the proposed cascade with nested dichotomies. Each layer enforces a semantically aligned gate. Samples flow top-down: at every stage, out-of-scope inputs are rejected and the remainder becomes progressively more specific, from open-world data to the task’s known world (ID).
@@ 4.2 Semantic OOD categories for MonuMAI
  • Monuments with known-styles: Finally, the known world or in-distribution (ID) set comprises the monuments belonging to the architectural styles used during the training of MonuMAI. This domain is explicitly defined by the training dataset employed in the development of the MonuMAI model.
@@ 4.3.1 Layer 0 - Far-world filtering
Based on the study by Zhang et al. using the OpenOOD benchmark [58], we selected the AdaScale-A postprocessor [42] detector for this layer, due to its excellent performance in far-OOD scenarios when ImageNet serves as the ID dataset. Moreover, the detector exhibits slightly reduced performance in near-OOD settings, meaning it may allow through samples visually similar to those in ImageNet when uncertainty arises. This behavior is not problematic at this stage, as it introduces a degree of tolerance that can later be refined by subsequent, more specialized filtering layers.
The detector was implemented using a RegNet-Y-16GF model pre-trained on ImageNet with SWAG augmentation as the backbone architecture. Following our definition of the MonuMAI near-world as the domain of color images representing real-world scenarios, we configured the system to use a subset of ImageNet images as the in-distribution reference, since ImageNet comprises numerous categories that also include monuments and buildings, thereby aligning with the near-world domain. For MonuMAI far-world representation, we constructed a mixed out-of-distribution dataset comprising MNIST handwritten digits and DTD texture patterns, which represent inputs that do not meet the conditions of real-world color photography scenarios and thus fall outside the near-world domain. The model extracts 3024-dimensional feature representations from the avgpool layer, and the AdaScale-A postprocessor applies adaptive scaling weights computed from class-wise variance statistics during the setup phase. The implementation employs standard ImageNet preprocessing with 384×384 pixel resolution and channel-wise normalization. The optimal decision threshold is determined through F1-score maximization on a validation subset.
@@ 4.3.4 Layer 3 - Known vs Unknown Style
This stage is formulated as an OOD detection problem over the MonuMAI model itself, which has already been trained on its corresponding ID dataset. Additionally, the monument dataset constructed for the previous layer can be leveraged to support outlier exposure, thereby facilitating the discrimination between known and unknown architectural styles.
As in previous layers, we rely on the OpenOOD benchmark to select a suitable detector. In this case, the setting corresponds to a near-OOD scenario, and the dataset is entirely custom-built, unlike in layer zero, where ImageNet was used as the ID distribution. For this reason, we selected entropy-based detection with outlier exposure fine-tuning  [21] [20], a detector that demonstrates strong performance in near-OOD detection across multiple ID datasets and supports training with outlier exposure.
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{ID}}+\lambda\times\mathcal{L}_{\text{OE}}$
where $\mathcal{L}_{\text{ID}}$ represents the standard cross-entropy loss computed on in-distribution samples (known architectural styles from the MonuTrain dataset), and $\mathcal{L}_{\text{OE}}$ implements the outlier exposure loss that encourages uniform probability distributions over unknown architectural styles. The outlier exposure loss is formulated as a KL divergence between the model’s predicted probabilities for out-of-distribution samples and a uniform distribution:
@@
For training, we utilize a custom dataset of monument images constructed through the Wikimedia Commons scraping procedures and CLIP-based thematic filtering described in previous sections (Layer 2), comprising architectural styles from across the globe that are not present in the MonuMAI ID training dataset. These globally diverse monument images serve as outlier samples, ensuring that the model learns to recognize architectural styles beyond those found in the Iberian Peninsula.
@@ 5.1 Experimental setup
In total, MonuTest-OOD contains 6354 images distributed across the different OOD and ID categories. The detailed category distribution is provided in Table 2.
@@
Known monuments (ID) 1038
Total (ID + OOD) 6354 (1038 + 5316)
@@
Additionally, a qualitative analysis will also be conducted on the images filtered at each layer of the system, with the aim of drawing relevant conclusions regarding the individual behavior of each layer. The OOD images misclassified as monuments will also be analized and compared between the SeNeDiF-OOD methodology and the MonuMAI original OOD pipeline.
@@ 5.3.1 Results by layers
AUROC Accuracy Precision Recall ID F1 Specificity OOD FPR95
@@
ID class: near-world images (pass candidates).
@@
Note: FPR95 computed at 95% TPR on ID. Precision and F1 are derived from (Accuracy, Recall, Specificity) under the test-set class prior.
@@
Table 5: Overall performance for layer 2 (monument vs building classification) of the SeNeDiF-OOD methodology by subcategory.
@@
ID class: monuments (known + unknown).
@@
ID class: monuments (known + unknown).
@@
ID class: known-style monuments.

### v3

@@ 1 Introduction
Drawing a parallel with multi-class classification, where each class would represent either an in-distribution (in-distribution) input or a type of OOD input, we can consider nested dichotomies techniques [15] to address this problem. Nested dichotomies break down classification tasks into multiple binary subproblems, each contrasting one category against the rest. During testing, each step determines whether the sample belongs to the split class. If it does not, the process continues to the next dichotomy. This sequence repeats until the dichotomy that assigns a class to the input is reached. Generally, nested dichotomies are less favored over other decomposition techniques like one-vs-one or one-vs-all [28] because, despite being more efficient, their performance depends significantly on the hierarchy of divisions within the dichotomies. However, when categories are already hierarchically structured, this limitation turns into an advantage.
@@
Many computer vision deployments still operationalize OOD recognition as a single scalar score that separates in-distribution from OOD under an essentially unstructured notion of novelty. This abstraction is misaligned with open-world, user-facing pipelines, where OOD inputs arise from heterogeneous mechanisms and demand different responses, ranging from unrelated content to semantically adjacent but unsupported classes and acquisition degradations. Consequently, a monolithic score offers limited interpretability and provides weak guidance for system-level handling. A missing ingredient is a principled methodology that decomposes OOD recognition into semantically meaningful decisions, enabling robust filtering while characterizing distinct OOD modes.
@@
A structured view of novelty can be obtained by organizing the input space into successive semantic checkpoints rather than treating it as a flat in-distribution versus OOD partition. In such a design, each checkpoint captures a coarse-to-fine distinction that mirrors how a practitioner would rule out implausible content before committing to a specific interpretation. This hierarchical organization encourages the use of specialized criteria at each stage and constrains error propagation, since ambiguous samples are intercepted early by broad semantic tests instead of being forced into a final decision. Beyond improving reliability, this methodology naturally yields a diagnostic outcome that indicates where and why an input becomes incompatible with the intended operating domain, which is valuable for monitoring, curation, and data acquisition strategies.
@@
In this paper, we introduce SeNeDiF-OOD (Semantic Nested Dichotomies Fusion for OOD detection), a new methodology that casts OOD recognition as a structured fusion process over a semantic cascade of dichotomous decisions. In particular, SeNeDiF-OOD combines hierarchical decision-level fusion with semantic priors embedded in the nested dichotomies topology. These priors provide human-aligned guidance that facilitates OOD detection. At each level of the hierarchy, the method also integrates heterogeneous expert fusion, leveraging the representational capabilities of artificial intelligence (AI) models to better separate the underlying dichotomies. Overall, this design enables a principled fusion of human knowledge and machine intelligence, which can be beneficial for facilitating OOD detection, a setting in which many computer vision models still struggle [61].
@@
Using a dataset composed of images uploaded by MonuMAI users, complemented with additional images obtained from public datasets covering different domains, we conducted an experiment to compare SeNeDiF-OOD methodology against the MonuMAI deep learning model itself. The results confirm that SeNeDiF-OOD significantly reduces false detections of monuments compared to the original model while preserving in-distribution images. Additionally, we analyze the system layer by layer, evaluating the effectiveness of each layer individually and drawing conclusions about the types of OOD images filtered at each stage.
@@
The structure of this paper is as follows. Section 2 provides the necessary background on out-of-distribution detection and nested dichotomies. Section 3 establishes the theoretical foundations of the proposed SeNeDiF-OOD methodology, formulating a generalizable methodology for hierarchical decision fusion. Section 4 details the practical implementation of the SeNeDiF-OOD methodology within the MonuMAI framework, outlining the established layers and the training procedures for each. Section 5 presents the experiments conducted on this dichotomy-based system, comparing its performance with the original MonuMAI version and analyzing the results both globally and at the layer level. Finally, Section 6 concludes with a summary of the work and key takeaways.
@@ 2.1 Out-of-distribution detection
Out-of-distribution detection has emerged as a fundamental component for ensuring the robustness, safety, and trustworthiness of modern machine learning systems [55, 18]. In real-world deployment, models frequently encounter inputs that deviate significantly from the statistical distribution of their training data. These inputs, referred to as OOD samples, can originate from different semantic categories, domains, acquisition conditions, or even adversarial manipulations. If not properly identified, OOD samples may be misclassified with high confidence, potentially leading to severe consequences in high-stakes domains such as autonomous driving, healthcare, or financial systems.
The importance of OOD detection lies in its role as a safeguard mechanism. For example, an autonomous driving system trained primarily on urban street scenes should raise an alert when faced with unexpected rural or adverse weather conditions, instead of producing overconfident predictions that may compromise safety [23, 39]. Similarly, in medical imaging, diagnostic systems must detect when a sample corresponds to a rare disease or a novel variant not represented in training data, in order to avoid erroneous diagnoses [27, 53, 5]. Likewise, in natural language processing, the rise of foundation models and their applicability as autonomous agents makes it necessary to further develop robust techniques for handling text inputs that fall outside the scope anticipated during system design [26, 56]. The primary goals of OOD detection are therefore to improve system reliability by preventing harmful misclassifications, to facilitate safer deployment of machine learning models, and to enable iterative model refinement by flagging unexpected cases for human inspection [4].
@@ 2.2 Nested dichotomies for multi-class problems
Nested dichotomies [15] constitute a structured decomposition framework for tackling multi-class classification problems by reducing them into a hierarchy of binary decisions. Unlike traditional strategies such as one-vs-one or one-vs-all, nested dichotomies organize class discrimination into a binary tree structure, where each internal node corresponds to a binary classifier that separates a subset of classes into two disjoint partitions. This recursive partitioning continues until each node corresponds to a single class.
@@
3 The SeNeDiF-OOD Methodology for Image Classification
@@
Section 3.1 introduces a taxonomy of OOD image types, illustrating the variety of OOD scenarios that can arise across different applications. Section 3.2 formalizes the SeNeDiF-OOD architecture, constructing a semantically grounded hierarchy that maps specific OOD categories to distinct decision fusion nodes. Lastly, Section 3.3 situates the proposed methodology within the broader context of responsible artificial intelligence, emphasizing key considerations such as robustness, explainability, and human alignment.
@@ 3.1 Open-world image classification: identifying OOD categories
  • Malicious OOD inputs. They are often stem from adversarial attacks or other exploits that leverage the model weaknesses. These images are not only out-of-distribution but intentionally crafted to deceive the model, exploiting weaknesses in the decision boundaries of the models. Unlike accidental OOD inputs, these are targeted and present an elevated risk in high-stakes applications.
@@
  • Degraded images. These include inputs with heavy noise, extreme lighting conditions, motion blur or compression artifacts. Although these images may still be of scope in terms of subject matter, the training distribution of the models is not likely to cover these distortions comprehensively.
@@ 3.2 SeNeDiF-OOD World Hierarchy and Layers
  • The known-world: the exact training distribution. Finally, the known-world represents the precise training distribution on which the model was trained on. The last layer has the goal of filtering out the novel classes that are not represented in the training set, but are considered to belong to the full-problem world. For the last layer, it will be usually necessary to resort to a specialized OOD detector, since it may not be feasible to have a full distribution of the problem unknown world. Detectors with outlier exposure [22] may help to define a clearer border between known and unknown worlds and facilitate novelty detection. This world represents the comfort zone of the model, where it can make confident predictions based on its learned parameters. This world is the in-distribution world: no OOD inputs are expected in this world.
This hierarchical structure of OOD categories can be naturally addressed using nested dichotomies, by directly extending the approach typically employed for multi-class problems. In this context, each OOD category can be treated as a distinct class, after excluding its internal categories to prevent overlap. Furthermore, the hierarchical organization of the categories enables the construction of a canonical and semantically meaningful dichotomy tree. At each node, the objective is to distinguish the outermost category from the remaining ones, progressively refining the classification until reaching the in-distribution category in the final split. This design avoids reliance on a randomly generated dichotomy tree, providing a more structured and interpretable decision process.
@@
Figure 3: Schematic of the proposed cascade with nested dichotomies. Each layer enforces a semantically aligned gate. Samples flow top-down: at every stage, out-of-scope inputs are rejected, and the remainder becomes progressively more specific, from open-world data to the task’s in-distribution data.
@@ 4.2 Semantic OOD Categories for MonuMAI
  • Monuments with known-styles: Finally, the known world or in-distribution set comprises the monuments belonging to the architectural styles used during the training of MonuMAI. This domain is explicitly defined by the training dataset employed in the development of the MonuMAI model.
@@ 4.3.1 Layer 0 - Far-world filtering
Based on the study by Zhang et al. using the OpenOOD benchmark [58], we selected the AdaScale-A postprocessor [42] detector for this layer, due to its excellent performance in far-OOD scenarios when ImageNet serves as the in-distribution dataset. Moreover, the detector exhibits slightly reduced performance in near-OOD settings, meaning it may allow through samples visually similar to those in ImageNet when uncertainty arises. This behavior is not problematic at this stage, as it introduces a degree of tolerance that can later be refined by subsequent, more specialized filtering layers.
The detector was implemented using a RegNet-Y-16GF model pre-trained on ImageNet with Stochastic Weight Averaging- Gaussian augmentation as the backbone architecture. Following our definition of the MonuMAI near-world as the domain of color images representing real-world scenarios, we configured the system to use a subset of ImageNet images as the in-distribution reference, since ImageNet comprises numerous categories that also include monuments and buildings, thereby aligning with the near-world domain. For MonuMAI far-world representation, we constructed a mixed out-of-distribution dataset comprising MNIST handwritten digits and DTD texture patterns, which represent inputs that do not meet the conditions of real-world color photography scenarios and thus fall outside the near-world domain. The model extracts 3024-dimensional feature representations from the avgpool layer, and the AdaScale-A postprocessor applies adaptive scaling weights computed from class-wise variance statistics during the setup phase. The implementation employs standard ImageNet preprocessing with 384×384 pixel resolution and channel-wise normalization. The optimal decision threshold is determined through F1-score maximization on a validation subset.
@@ 4.3.4 Layer 3 - Known vs Unknown Style
This stage is formulated as an OOD detection problem over the MonuMAI model itself, which has already been trained on its corresponding in-distribution dataset. Additionally, the monument dataset constructed for the previous layer can be leveraged to support outlier exposure, thereby facilitating the discrimination between known and unknown architectural styles.
As in previous layers, we rely on the OpenOOD benchmark to select a suitable detector. In this case, the setting corresponds to a near-OOD scenario, and the dataset is entirely custom-built, unlike in layer zero, where ImageNet was used as the in-distribution class. For this reason, we selected entropy-based detection with outlier exposure fine-tuning  [21] [20], a detector that demonstrates strong performance in near-OOD detection across multiple in-distribution datasets and supports training with outlier exposure.
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{in-distribution}}+\lambda\times\mathcal{L}_{\text{OE}}$
where $\mathcal{L}_{\text{in-distribution}}$ represents the standard cross-entropy loss computed on in-distribution samples (known architectural styles from the MonuTrain dataset), and $\mathcal{L}_{\text{OE}}$ implements the outlier exposure loss that encourages uniform probability distributions over unknown architectural styles. The outlier exposure loss is formulated as a KL divergence between the model’s predicted probabilities for out-of-distribution samples and a uniform distribution:
@@
For training, we utilize a custom dataset of monument images constructed through the Wikimedia Commons scraping procedures and CLIP-based thematic filtering described in previous sections (Layer 2), comprising architectural styles from across the globe that are not present in the MonuMAI in-distribution training dataset. These globally diverse monument images serve as outlier samples, ensuring that the model learns to recognize architectural styles beyond those found in the Iberian Peninsula.
@@ 5.1 Experimental setup
In total, MonuTest-OOD contains 6354 images distributed across the different OOD and in-distribution categories. The detailed category distribution is provided in Table 2.
@@
Known monuments (in-distribution) 1038
Total (in-distribution + OOD) 6354 (1038 + 5316)
@@
Additionally, a qualitative analysis will also be conducted on the images filtered at each layer of the system, with the aim of drawing relevant conclusions regarding the individual behavior of each layer. The OOD images misclassified as monuments will also be analyzed and compared between the SeNeDiF-OOD methodology and the MonuMAI original OOD pipeline.
@@ 5.3.1 Results by layers
AUROC Accuracy Precision Recall (in-distribution) F1 Specificity OOD FPR95
@@
In-Distribution class: near-world images (pass candidates).
@@
Note: FPR95 computed at 95% TPR on in-distribution class. Precision and F1 are derived from (Accuracy, Recall, Specificity) under the test-set class prior.
@@
Table 5: Overall performance for layer 2 (monument vs. building classification) of the SeNeDiF-OOD methodology by subcategory.
@@
In-Distribution class: monuments (known + unknown).
@@
In-Distribution class: monuments (known + unknown).
@@
In-Distribution class: known-style monuments.